Two 2000 ?kg cars are traveling 95 km/h in opposite directions when they collide and are brought to rest.

Part A

Estimate the change in entropy of the universe as a result of this collision. Assume T=25?C.

Express your answer to two significant figures and include the appropriate units.
Answer #1

total energy of the two cars are = sum of their kinetic energy

90 km/h = 95*1000/3600 = 26.4 m/s
so total kinetic energy = {(2*0.5*2000*26.4 2)} = 1393920 j

change in entropy = (change in heat energy / temperature )

as body comes in the rest so no other form of energy are changing. only kinetic energy are converted into the heat loss.

hence, entropy change = (1393920/(25+273)) (chnage temperature in kelvin)

4677.58 j/K